OpenStudy (anonymous):

Write the first trigonometric function in terms of the second for θ in the given quadrant. csc θ, cot θ; θ in quadrant III

OpenStudy (anonymous):

missing something?

OpenStudy (anonymous):

nope thats the problem loll any idea?

OpenStudy (anonymous):

you have to find csc θ =

OpenStudy (anonymous):

yeah i see what it is asking

OpenStudy (anonymous):

you know \[\sin^2(\theta)+\cos^2(\theta)=1\] divide by \(\sin^2(\theta)\) to get \[1+\cot^2(\theta)=\csc^2(\theta)\]

OpenStudy (anonymous):

that means \[\csc(\theta)=\pm\sqrt{1+\cot^2(\theta)}\]

OpenStudy (anonymous):

in quadrant 3 cosecant is negative, so pick \[\csc(\theta)=-\sqrt{1+\cot^2(\theta)}\]

OpenStudy (anonymous):

yeah i was thinking the same but they gave csc(sine) but also cot which is tan

OpenStudy (anonymous):

can you please explain how you got 1+cot2(θ)=csc2(θ) when you divided by sin^2theta?

OpenStudy (e.mccormick):

\[\frac{\sin^2\theta}{\sin^2\theta}+\frac{\cos^2\theta}{\sin^2\theta}=\frac{1}{\sin^2 \theta}\]Now simplify that

OpenStudy (e.mccormick):

The Pythagorean Identities are the original and the two you can easily get. One of then you get by dividing through by sine, the other you get by dividing through by cosine.

OpenStudy (anonymous):